The product of two consecutive even numbers decreased by three times the larger number is fifty find BOTH numbers show all work
Answer by CubeyThePenguin(3113)

consecutive even numbers: x, (x+2)

(x)(x+2) - 3(x+2) = 50
x^2 + 2x - 3x - 6 = 50
x^2 - x - 56 = 0
(x - 8)(x + 7) = 0

x is even, so x = 8. The numbers are 8 and 10.